The pressure of the helium in the balloon is 100 000 Pa. The volume of the balloon is 0.030 m3 . The balloon is compressed at a constant temperature causing the volume to decrease to 0.025 m3 . No helium leaves the balloon. Calculate the new pressure in the balloon


1
Expert's answer
2021-12-28T08:44:07-0500

Apply the combined gas law keeping in mind the temperature did not change:


"\dfrac{PV}T=\dfrac{P_nV_n}T,\\\space\\
PV=P_nV_n,\\\space\\
P_n=\frac{PV}{V_n}=\dfrac{100000·0.03}{0.025}=120000\text{ Pa}."